OpenStudy (anonymous):

Find the equation of an ellipse with vertices (3, 0) and (-3, 0) and foci of (2, 0) and (-2, 0). I dont know which formula to use!!!

OpenStudy (anonymous):

Should i use the Major Axis Horizontal Formula? \[\frac{ (x-h)^{2} }{ a ^{2} }+\frac{ (y-k)^{2} }{ b ^{2} }\]

OpenStudy (anonymous):

Or the Major Axis Vertical Formula? \[\frac{ (x-h)^{2} }{ b ^{2} }+\frac{ (y-k)^{2} }{ a ^{2}}\]

OpenStudy (anonymous):

How do you know when to use which formula? Please I really need help on this!